OpenStudy (anonymous):

Solve |2x-1| + 3 = |5-x|

OpenStudy (anonymous):

2x + 2 = 5 - x 3x = 3 x = 1 2x + 2 = x - 5 x = -3

OpenStudy (anonymous):

after the first step, the -1 in the first absolute value wouldnt be subtraction anymore, so it wouldnt be +2, it would be +4, right?

OpenStudy (anonymous):

i think ive figured it out, i got .33 repeating or 1/3*

OpenStudy (anonymous):

i don't think that's right...

OpenStudy (anonymous):

you could also do: 4 - 2x = 5 - x x = -1 4 - 2x = x - 5 x = 3 so you should have four answers in total: 1, -1, 3, -3